OpenStudy (anonymous):

Hi I am having trouble getting this a/20+4/15=9/15

OpenStudy (anonymous):

\[\frac{ a }{ 20 } + \frac{ 4 }{ 15 } = \frac{ 9 }{ 15 }\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

adding fractions needs a common denominator. or, another better alternative is to multiply the entire equation, left and right side, by 20. This will cancel the denominator under the "a" and help get you ready to solve for a.

OpenStudy (anonymous):

20(a/20) + 20(4/15) = 20(9/15) Then simplify...

OpenStudy (anonymous):

Oh ok and then I'm done?

OpenStudy (anonymous):

yes, if you simplified, you hopefully got a + 80/15 = 180/15 and then... a = 100/15, which reduces to... a = 20/3

OpenStudy (anonymous):

wait, I think I made an error...

OpenStudy (anonymous):

Ohh ok thank you soo much!! :)

OpenStudy (anonymous):

no, it turns out a = 20/3 was correct :) Do you understand how to solve for the variable when there are fractions involved? Hope this helped :)